CathFoods will release a new range of candies which contain antioxidants. New equipment to manufacture the candy will cost

$ 3

​million, which will be depreciated by​ straight-line depreciation over

five

years. In​ addition, there will be​ $5 million spent on promoting the new candy line. It is expected that the range of candies will bring in revenues of

$ 5

million per year for

fivefive

years with production and support costs of​ $1.5 million per year. If​ CathFood's marginal tax rate is​ 35%, what are the incremental free cash flows in the second year of this​ project?

A.

$1.015

million

B.

$2.485

million

C.

$2.900

million

D.

$1.750

million
Answer #1

Promotion cost of $5 million should be ingnored as it is a sunk cost. The cost that cannot be recovered even if the project is abandoned.

Annual depreciation = 3 / 5 = 0.6

Incremental cash flow = (Revenue - costs - depreciation)(1 - tax) + depreciation

Incremental cash flow = (5 - 1.5 - 0.6)(1 - 0.35) + 0.6

Incremental cash flow = 1.885 + 0.6

Incremental cash flow = $2.485 million